# Generic torsion-free groups and Rubin actions

### v1

@@ Abstract.
We use model theoretic forcing to prove that a generic countable torsion-free group does not admit any nontrivial locally moving action on a Hausdorff topological space, and yet admits a rich Rubin poset.
@@ 1. Introduction
In this article, we investigate the question of whether a generic countable torsion-free group admits sufficiently rich actions on topological spaces. We are motivated primarily by the problem of deciding whether or not there exists a countable torsion-free group which admits no nontrivial action on a compact manifold.
@@
    Locally moving if for all nonempty $U\subseteq X$, we have $G_{U}$ is nontrivial.
@@
Such an action will be called a Rubin action. We will say that a group $G$ is a Rubin group if $G$ admits a Rubin action, and a weakly Rubin group if it admits a homomorphism to $\operatorname{Homeo}(X)$ for some Hausdorff space $X$ with at least two points, whose image is a locally moving group of homeomorphisms.
Let $G\leq\operatorname{Homeo}(X)$ be a Rubin group. The driving force behind Rubin’s Theorem (Theorem 1.1 above) is that from the local moving condition, one can recover a substantial amount of the topology of $X$. Specifically, one can recover a dense subset of the Boolean algebra $\operatorname{RO}(X)$ of regular open sets of $X$; this fact has been used to investigate the model theory of homeomorphism groups of manifolds; see [5, 6, 7]. Here, we say an open set is regular if it is equal to the interior of its closure. Regular open sets in $X$ arise from supports of homeomorphisms of $X$.
A homeomorphism $g$ of $X$ has an open support $\operatorname{supp}g$, which consists of the points in $X$ which are not fixed by $g$. We define the extended support of $g$, or $\operatorname{supp^{e}}g$ as the interior of the closure of $\operatorname{supp}g$; this is the smallest regular open set containing $\operatorname{supp}g$. We say that $g\in G$ is algebraically disjoint from $f\in G$ if for all $h\in G$ such that $[f,h]\neq 1$, there are elements $a,b\in C_{G}(g)$ such that
@@
Here, $C_{G}(g)$ is the centralizer of $g$ in $G$; note that algebraic disjointness makes sense in an arbitrary group and not just in a homeomorphism group. Also, note that if $g$ is algebraically disjoint from $f$, then $[g,f]=1$, for indeed choosing $h=g$ provides a contradiction. Algebraic disjointness is not necessarily a symmetric relation; for example, if $G$ is the symmetric group on $\{1,2,3,4\}$, the permutation (1 2)12(1\>2)( 1 2 ) is algebraically disjoint from (1 2)(3 4)1234(1\>2)(3\>4)( 1 2 ) ( 3 4 ), but not vice-versa.
@@
and let $C_{G}(S_{f})$ be the centralizer of $S_{f}$ in $G$. The group $C_{G}(S_{f})$ is nontrivial since it follows from the definition of algebraic disjointness that $f\in C_{G}(S_{f})$. Also, note that it is possible that $C_{G}(S_{f})=C_{G}(S_{f^{\prime}})$ for distinct elements $f,f^{\prime}\in G$. Let $\mathcal{P}(G)$ be the poset consisting of elements
[…]
@@
The main result of this paper is that “most” countable groups are not Rubin groups (or even weakly Rubin groups), yet admit a rich Rubin poset. Here, countable groups are organized into a Polish space of marked countable groups, which can then be investigated from a descriptive set theory point of view; see  [2]. In particular, the notions of meagerness and comeagerness make sense for the space of marked countable groups, and a property of groups is called generic if it holds for groups in a comeager subset of the space of marked countable groups; we will give a precise definitions in Section 2. The main result of this article is:
@@
    Two nontrivial elements of $G$ commute if and only if they are algebraically disjoint in $G$.
@@
We also remark that a generic torsion-free group has a rich subgroup structure. In fact it contains every finitely generated torsion-free group with a solvable word problem as a subgroup. In turn, the poset of cyclic subgroups ordered by inclusion is very rich: for instance, a generic torsion-free group will contain a copy of $\mathbb{Q}*_{\mathbb{Z}}\mathbb{Q}$, the amalgamated product of two copies of $\mathbb{Q}$ over their respective copies of the integers. We thus obtain many cyclic subgroups which are all distinct, but which contain a common cyclic subgroup. One can repeat a process of this amalgamation and taking direct sums of groups ad infinitum, thus building a small piece of the very complicated poset of cyclic subgroups of a generic torsion-free group. Also, we remark that any comeager subset of the space of torsion-free groups will contain continuum many groups up to isomorphism (this is a direct consequence of Theorem $1.1.6$ of [2]).
Theorem 1.2 is established through model theoretic forcing; the key is to find a Banach–Mazur game in which one player can force any pair of commuting elements which do not share a common power to be mutually algebraically disjoint. The same player also forces any two nontrivial elements to be conjugate, resulting in a simple compiled group. We also emphasize the fact that our result is much easier to prove if the torsion-free requirement is dropped, though the requirement that the group be torsion-free is central to our program of investigating obstructions to group actions on compact manifolds: indeed, the existence of torsion is often an elementary source of such obstructions, and should be excluded for the development of a deeper theory.
@@ 2.1. Algebraic disjointness
    If $f,g\in\Gamma$ satisfy $(\operatorname{supp}g)\cap(\operatorname{supp}f)=\varnothing$ then $f$ and $g$ are algebraically disjoint.
@@
    If $f,g\in\Gamma$ are algebraically disjoint then $(\operatorname{supp}g^{12})\cap(\operatorname{supp}f)=\varnothing$.
@@ 2.2. The space of enumerated groups
Let $\underline{x}$ denote a finite tuple of variables. We let $\Sigma(\underline{x})$ be a quantifier-free formula in the language of group theory consisting of finitely many equations and inequations of the form $w(\underline{x})=1$ or $w(\underline{x})\neq 1$. Such a $\Sigma(\underline{x})$ is called a system. If $\underline{g}$ is a finite tuple of natural numbers, the system $\Sigma(\underline{g})$ defines a clopen set $U_{\Sigma}$ in $\mathcal{G}$ by considering the groups in which $\Sigma(\underline{g})$ holds, and these clopen sets determine a basis for the topology of $\mathcal{G}$.
@@ 2.3. Banach–Mazur games
Let $P$ be a property of countable groups, and let $\mathcal{G}_{P}$ be the subset of $\mathcal{G}$ consisting of groups satisfying $P$. A Banach–Mazur game [3, 9] is a game in which players A and B take turns choosing finite systems $\{\Sigma_{i}(\underline{g})\}_{i\in\mathbb{N}}$ such that the clopen sets $U_{i}=U_{\Sigma_{i}}$ are nested and whose intersection determines a unique group in $\mathcal{G}$ called the compiled group. Player A wins if the compiled group lies in $\mathcal{G}_{P}$. Player A has a winning strategy if for any position $U_{n}$ in the game, there is a possible choice of systems for which player A wins. It is a standard fact that player A has a winning strategy if and only if $\mathcal{G}_{P}$ is comeager, which is to say that $\mathcal{G}_{P}$ contains a countable intersection of dense open sets. We say that $P$ is a generic property of groups if this holds. Note that a countably infinite conjunction of generic properties is generic, from the Baire category theorem.
@@ 3. Some combinatorial group theory
We will always assume that abstract relations in a group are reduced and cyclically reduced. Moreover, when dealing with a group presentation, we will always assume that the set of relations is closed under taking inverses and cyclic permutations. Our standing notation will be that $G=\langle g,h\mid\mathcal{R}_{0}\rangle$ is a subgroup of larger ambient torsion-free group $\langle f,g,h\rangle$, wherein $[f,h]\neq 1$ but $[f,g]=1$. Note that this implies that $h\notin\langle g\rangle$ and that if $G$ is cyclic then $g=h^{n}$ for some $n\in\mathbb{Z}$ with $|n|>1$.
@@
Let $G$ be nontrivial and torsion-free group. Suppose furthermore that $g,h\neq 1$, and if $G$ is cyclic then $g=h^{n}$ for some $n\in\mathbb{Z}$ with $|n|>1$. Then there exists a finitely generated torsion-free group $\Gamma$ such that:
@@
    There exists an $a,b\in\Gamma\setminus 1$ centralizing $g\in G$.
@@
    The commutator $[a,[b,h]]\neq 1$ and centralizes $g$.
To prove Lemma 3.1 for the non-cyclic case, we need to first establish the following lemma.
@@
Let $K=\langle g,\gamma\rangle$ be a nontrivial and torsion-free group, and suppose that for all nonzero $n\in\mathbb{Z}$ we have $\gamma^{n}\notin\langle\langle g\rangle\rangle_{K}$. Then there is a finitely generated torsion-free overgroup $\Gamma$ and a nontrivial $a\in\Gamma$ such that:
@@
    The commutator $[a,\gamma]$ is nontrivial and centralizes $g$.
@@
Let $K_{1}$ and $K_{2}$ be two isomorphic copies of $K$, with $g_{i},\gamma_{i}\in G_{i},1\leq i\leq 2$ being the corresponding generators in the two copies. Set
@@
Observe that $G_{3}$ is isomorphic to $G$ via the map $\phi$ sending
@@
Indeed, if the map $\phi$ is a well-defined homomorphism of groups then it is clearly surjective; moreover, by projection onto the first factor, we see that it must be injective as well. Thus, it suffices to show that if $w(g,\gamma)$ is a relation in $g$ and $\gamma$ then $\phi(w(g,\gamma))$ is the identity. By the definition of $\phi$, we see that $\phi(w(g,\gamma))$ will be the identity if and only if $w$ has zero exponent sum in $\gamma$. Since we assumed that $\gamma^{n}\notin\langle\langle g\rangle\rangle_{G}$ for all nonzero $n$, it is immediate that any relation in $G$ will have zero exponent sum in $\gamma$.
It thus suffices to find the desired element $a\in C_{\Gamma}(g)$; we simply take a generator $a\in\mathbb{Z}$ of the infinite cyclic group in the free product. By construction, $(1,a)$ commutes with $(g_{1},1)\in G_{3}$, and the commutator $[(1,a),(\gamma_{1},\gamma_{2})]$ is nontrivial and commutes with $(g_{1},1)$. ∎
@@
Let $G=\langle g,h\rangle$ be a non-cylic and torsion-free group, and let $H$ be a free product with amalgamation of two copies $G_{1}=\langle g_{1},h_{1}\rangle$ and $G_{2}=\langle g_{2},h_{2}\rangle$ of $G$, wherein the subgroups $\langle g_{1}\rangle$ and $\langle g_{2}\rangle$ being amalgamated into a single cyclic group $\langle g\rangle$. Let $K=\langle h_{2}^{-1}h_{1},g\rangle$. Then for all nonzero $n\in\mathbb{Z}$, we have
@@
In the case where $G$ is cyclic, then it must be the case that for some $n\in\mathbb{Z}$ with $|n|>1$ we have that $h^{n}=g$. Indeed, if $g^{n}=h$ then $h$ must commute with $f$ which contradicts our assumption. We may produce the required overgroup by amalgamating a copy of $\mathbb{Z}^{3}$ generated by elements $\{a,b,c\}$ with $G$, by identifying the cyclic subgroups $\langle c\rangle$ and $\langle g\rangle$. Using the normal form for amalgamated free products, it is clear that $[a,[b,h]]\neq 1$. Moreover,
@@
since $[\alpha,g]=1$ for each $\alpha\in\{a,b,h\}$.
Assume now that $G$ is noncyclic. Let $Q$ be the group $H$ obtained by amalgamating two copies of $G$ along their respective copies of $g$, and let $\beta$ be the automorphism of $Q$ which fixes $g$ and exchanges the two copies $h_{1}$ and $h_{2}$ of $h$ in $Q$, coming from the two amalgamated subgroups. We set $\Gamma_{0}$ to be the semidirect product of $Q$ with $\mathbb{Z}$, where $\mathbb{Z}$ acts on $Q$ by $\beta$. Clearly $\Gamma_{0}$ is torsion-free.
We identify the generator of $\mathbb{Z}$ as an element $b$ of $\Gamma_{0}$, and we note that $b$ centralizes $g$. Observe that $[b,h_{1}]=h_{2}^{-1}h_{1}$; we call this commutator $\gamma$, which clearly represents a nontrivial element of $\Gamma_{0}$.
Let $K=\langle g,\gamma\rangle$. By Lemma 3.3, we have that $\gamma^{n}\notin\langle\langle g\rangle\rangle_{K}$ for any $n\in\mathbb{Z}\setminus\{0\}$, and so by Lemma 3.2, the group $K$ embeds in a torsion-free overgroup $\tilde{K}$, wherein some nontrivial $a\in\tilde{K}$ centralizes $g$ and satisfies $[a,\gamma]\neq 1$ and $[a,\gamma]$ centralizes $g$. Now, amalgamate $\tilde{K}$ and $\Gamma_{0}$ over their respective copies of $K$, and call the resulting group $\Gamma$. In $\Gamma$, we have
@@
and all three of $a$ and $b$ and $[a,[b,h_{1}]]$ centralize $g$. Since $\langle g,h_{1}\rangle$ is an isomorphic copy of $G$, we are done. ∎
We now need only prove Lemma 3.3. For this, we use Van Kampen diagrams over $H$; of course, Van Kampen diagrams will generally depend on a particular choice of presentation. When we construct $H$, we have two copies of $G$, and as such there are two copies of the set of relations $\mathcal{R}_{0}$, which we call $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, respectively. Thus, relations in $\mathcal{R}_{i}$ involve $g=g_{1}=g_{2}$ and $h_{i}$ only. We fix the resulting presentation for $H$ once and for all.
@@
Suppose that in $G$ there are nonzero integers $n,m\in\mathbb{Z}$ such that $hg^{n}h^{-1}=g^{m}$. Then in the group $H$, we have $[g^{n},h_{2}^{-1}h_{1}]=1$.

### v2

@@ Abstract.
We use model theoretic forcing to prove that a generic countable torsion-free group does not admit any non-trivial locally moving action on a Hausdorff topological space, and yet admits a rich Rubin poset.
@@ 1. Introduction
In this article, we investigate the question of whether a generic countable torsion-free group admits sufficiently rich actions on topological spaces. We are motivated primarily by the problem of deciding whether or not there exists a countable torsion-free group which admits no non-trivial action on a compact manifold; for instance, countable groups acting faithfully on connected compact 11–manifolds are characterized by linear or circular orderability. For manifolds of dimension two or higher, there are no known obstructions for a torsion-free countable group to act faithfully by homeomorphisms, though it seems likely that there are countable torsion-free groups which admit no non-trivial action by homeomorphisms on any compact manifold.
@@
    Locally moving if for all nonempty $U\subseteq X$, we have $G_{U}$ is non-trivial.
@@
A locally dense faithful action of a group on a locally compact Hausdorff space $X$ with no isolated points will be called a Rubin action. We will say that a group $G$ is a Rubin group if $G$ admits a Rubin action, and a weakly Rubin group if it admits a homomorphism to $\operatorname{Homeo}(X)$ for some Hausdorff space $X$ with at least two points, whose image is a locally moving group of homeomorphisms.
Let $G\leq\operatorname{Homeo}(X)$ be a Rubin group. The driving force behind Rubin’s Theorem (Theorem 1.1 above) is that from the local density condition, one can recover a substantial amount of the topology of $X$. Specifically, one can recover a dense subset of the Boolean algebra $\operatorname{RO}(X)$ of regular open sets of $X$; here, a subset $D$ of $X$ is dense if for every nonempty $U\in X$, there is a nonempty $V\in D$ such that $V\subseteq U$. This fact has been used to investigate the model theory of homeomorphism groups of manifolds; see [6, 7, 8]. Here, we say an open set is regular if it is equal to the interior of its closure. Regular open sets in $X$ arise from supports of homeomorphisms of $X$.
A homeomorphism $g$ of $X$ has an open support $\operatorname{supp}g$, which consists of the points in $X$ which are not fixed by $g$. We define the extended support of $g$, or $\operatorname{supp^{e}}g$ as the interior of the closure of $\operatorname{supp}g$; this is the smallest regular open set containing $\operatorname{supp}g$. We say that $g\in G$ is algebraically disjoint from $f\in G$ if for all $h\in G$ such that $f,h]\neq 1$, there are elements $a,b\in C_{G}(g)$ such that
@@
Here, $C_{G}(g)$ is the centralizer of $g$ in $G$; note that algebraic disjointness makes sense in an arbitrary group and not just in a homeomorphism group. Also, note that if $g$ is algebraically disjoint from $f$, then $[g,f]=1$, for indeed choosing $h=g$ provides a contradiction. It is a trivial though useful observation that in an abelian group, two elements are algebraically disjoint if and only if they commute, since the quantification over $h$ not commuting with $f$ is vacuous.
Algebraic disjointness is not necessarily a symmetric relation; for example, if $G$ is the symmetric group on $\{1,2,3,4\}$, the permutation (1 2)(1\>2) is algebraically disjoint from (1 2)(3 4)(1\>2)(3\>4), but not vice-versa.
@@
and let $C_{G}(S_{f})$ be the centralizer of $S_{f}$ in $G$. The group $C_{G}(S_{f})$ is non-trivial since it follows from the definition of algebraic disjointness that $f\in C_{G}(S_{f})$. Also, note that it is possible that $C_{G}(S_{f})=C_{G}(S_{f^{\prime}})$ for distinct elements $f,f^{\prime}\in G$. Let $\mathcal{P}(G)$ be the poset consisting of elements
$\{C_{G}(S_{f_{1}})\cap\cdots\cap C_{G}(S_{f_{n}})\mid f_{1},\ldots,f_{n}\in G,\,1\leq n\in\mathbb{N}\},$
@@
The main result of this paper is that “most” countable groups are not Rubin groups (or even weakly Rubin groups), yet admit a rich Rubin poset. Here, countable groups are organized into a Polish space of marked countable groups, which can then be investigated from a descriptive set theory point of view; see  [2]. In particular, the notions of meagerness and comeagerness make sense for the space of marked countable groups, and a property of groups is called generic if it holds for groups in a comeager subset of the space of marked countable groups. Properties of generic groups are studied by certain (infinite length) two-player games, and the class of groups obtained at the “end” of the game is called the class of compiled groups; we will give a precise definitions in Section 2. The main result of this article is:
@@
    Two non-trivial elements of $G$ commute if and only if they are algebraically disjoint in $G$.
@@
Remark 1.3.
A generic torsion-free group has a rich subgroup structure. In fact, it contains every finitely generated torsion-free group with a solvable word problem as a subgroup. In turn, the poset of cyclic subgroups ordered by inclusion is very rich: for instance, a generic torsion-free group will contain a copy of $\mathbb{Q}*_{\mathbb{Z}}\mathbb{Q}$, the amalgamated product of two copies of $\mathbb{Q}$ over their respective copies of the integers. We thus obtain many cyclic subgroups which are all distinct, but which contain a common cyclic subgroup. One can repeat a process of this amalgamation and taking direct sums of groups ad infinitum, thus building a small piece of the very complicated poset of cyclic subgroups of a generic torsion-free group. Also, we remark that any comeager subset of the space of torsion-free groups will contain continuum many groups up to isomorphism (this is a direct consequence of Theorem $1.1.6$ of [2]).
Theorem 1.2 is established through model theoretic forcing; the key is to find a Banach–Mazur game in which one player can force any pair of commuting elements which do not share a common power to be mutually algebraically disjoint. The same player also forces any two non-trivial elements to be conjugate, resulting in simple compiled groups. We also emphasize the fact that our result is much easier to prove if the torsion-free requirement is dropped, though the requirement that the group be torsion-free is central to our program of investigating obstructions to group actions on compact manifolds: indeed, the existence of torsion is often an elementary source of such obstructions, and should be excluded for the development of a deeper theory.
@@ 2.1. Algebraic disjointness
    If $f,g\in\Gamma$ satisfy $(\operatorname{supp}g)\cap(\operatorname{supp}f)=\varnothing$ then $f$ and $g$ are mutually algebraically disjoint.
@@
    If $f,g\in\Gamma$ are such that $g$ is algebraically disjoint from $f$, then $(\operatorname{supp}g^{12})\cap(\operatorname{supp}f)=\varnothing$.
@@ 2.2. The space of enumerated groups
Let $\underline{x}$ denote a finite tuple of variables. We let $\Sigma(\underline{x})$ be a quantifier-free formula in the language of group theory consisting of finitely many equations and inequations of the form $w(\underline{x})=1$ or $w(\underline{x})\neq 1$. Such a $\Sigma(\underline{x})$ is called a finite system. If $\underline{g}$ is a finite tuple of natural numbers, the system $\Sigma(\underline{g})$ defines a clopen set $U_{\Sigma}$ in $\mathcal{G}$ by considering the groups in which $\Sigma(\underline{g})$ holds, and these clopen sets determine a basis for the topology of $\mathcal{G}$.
@@ 2.3. Banach–Mazur games
Let $P$ be a property of countable groups, and let $\mathcal{G}_{P}$ be the subset of $\mathcal{G}$ consisting of groups satisfying $P$. A Banach–Mazur game [3, 10] is a game in which players A and B take turns choosing finite systems $\{\Sigma_{i}(\underline{g})\}_{i\in\mathbb{N}}$ such that the nonempty clopen sets $U_{i}=U_{\Sigma_{i}}$ are nested and whose intersection determines a class of groups in $\mathcal{G}$ called the compiled groups of the game. The nestedness of these open sets is crucial, since it prevents players from placing inconsistent conditions on the compiled groups. Any group $G\in\bigcap_{i}U_{i}$ is called a compiled group for the game. One should imagine that $A$ and $B$ are building a (class of groups) together, by gradually naming the elements of the group and expressing, finitely many equations and inequations at a time, the (partial) multiplication table of a (class of) groups. It may be the case that the full multiplication table can be determined over the course of a game, in which case the compiled group is unique. The nonemptiness of the sets $\{U_{i}\}_{i\in\mathbb{N}}$ is simply expressing the consistency of the finite systems played up to that round of the game; in particular, at least one group satisfies the conditions forced at any round.
At each stage, both $A$ and $B$ may only play a finite system (and in particular not a family of finite systems). On any turn, either player may play (any subset of) the finite systems that have been played up to that round; this amounts to that player wasting a turn. For the sake of convenience, we assume that after $n$ rounds, all natural numbers below $2n$ have been played (and possibly more).
Player A wins if the compiled groups lie in $\mathcal{G}_{P}$. Player A has a winning strategy if for any possible position $U_{n}$ in the game, there is a possible choice of systems for which player A wins, regardless of how player B plays; the choice of systems played by player A need not be fixed beforehand, but can depend on exactly how player B proceeds. It is a standard fact that player A has a winning strategy if and only if $\mathcal{G}_{P}$ is comeager, which is to say that $\mathcal{G}_{P}$ contains a countable intersection of dense open sets. We say that $P$ is a generic property of groups if this holds. Note that a countably infinite conjunction of generic properties (i.e. simultaneously requiring all of the countably many properties) is generic, from the Baire category theorem; indeed, the set of groups with each of the generic properties contains a countable intersection of dense open sets, and so the set of groups with countably many prescribed generic properties also contains a countable intersection of dense open sets.
Similarly, player B wins if the compiled groups lie in the complement of $\mathcal{G}_{P}$, and has a winning strategy if for any position $U_{n}$ in the game, there is a possible choice of systems for which player B wins. For instance, suppose $P$ is the property of having at least two conjugacy classes (in a torsion-free group). Without loss of generality, 0 has been declared to be the identity by player A in the first round. At every turn, player B chooses an $n\in\mathbb{N}$ exceeding the maximum of the numbers that have been played by either player, declares $n$ to be distinct from $\{0,\ldots,n-1\}$, and declares that $n$ conjugates $k$ to $k+1$ for $k\geq 1$. This is easily seen to be a winning strategy for player B.
2.4. Divisible groups
In later parts of this paper, divisible groups will arise. A general reference on divisible groups is [4]. Divisible groups are usually (though not always, like in this paper) abelian groups $A$ wherein for all $x\in A$, the equation $n\cdot y=x$ has a solution $y\in A$. There is an obvious generalization to nonabelian groups, wherein one requires solutions to the equation $y^{n}=x$. To allow for torsion-free divisible groups, one need not require $y$ to be non-trivial.
Abelian divisible groups have a well-defined torsion subgroup. An abelian torsion-free divisible group is isomorphic to a Cartesian power of the rationals, say $\mathbb{Q}^{I}$. The cardinality of the set $I$ is called the rank of the divisible group.
@@ 3. Some combinatorial group theory
We will always assume that abstract relations in a group are reduced and cyclically reduced. Moreover, when dealing with a group presentation, we will always assume that the set of relations is closed under taking inverses and cyclic permutations. Our standing notation will be that $G=\langle g,h\mid\mathcal{R}_{0}\rangle$ is a subgroup of larger ambient torsion-free group $\langle f,g,h\rangle$, wherein $f,h]\neq 1$ but $[f,g]=1$. Note that this implies that $h\notin\langle g\rangle$ and that if $G$ is cyclic then $g=h^{n}$ for some $n\in\mathbb{Z}$ with $|n|>1$.
@@
Let the group $G$ defined above be non-trivial and torsion-free group. Suppose furthermore that $g,h\neq 1$, and if $G$ is cyclic then $g=h^{n}$ for some $n\in\mathbb{Z}$ with $|n|>1$. Then there exists a finitely generated torsion-free group $\Gamma$ such that:
@@
    There exist distinct elements $a,b\in\Gamma\setminus 1$ centralizing $g\in G$.
@@
    The commutator $a,[b,h]]\neq 1$ and centralizes $g$.
We will prove Lemma 3.1 in both the cyclic and non-cyclic cases below within the body of the proof, though in order to prove Lemma 3.1 in the non-cyclic case, we will need to establish the following lemma.
@@
Let $K=\langle g,\gamma\rangle$ be a non-trivial and torsion-free group, and suppose that for all nonzero $n\in\mathbb{Z}$ we have $\gamma^{n}\notin\langle\langle g\rangle\rangle_{K}$. Then there is a finitely generated torsion-free overgroup $\Gamma$ and a non-trivial $a\in\Gamma$ such that:
@@
    The commutator $a,\gamma]$ is non-trivial and centralizes $g$.
@@
Let $K_{1}$ and $K_{2}$ be two isomorphic copies of $K$, with $g_{i},\gamma_{i}\in K_{i},1\leq i\leq 2$ being the corresponding generators in the two copies. Set
@@
Observe that $K_{3}$ is isomorphic to $K$ via the map $\phi$ sending
@@
Indeed, if the map $\phi$ is a well-defined homomorphism of groups then it is clearly surjective; moreover, by projection onto the first factor, we see that it must be injective as well. Thus, it suffices to show that if $w(g,\gamma)$ is a relation in $g$ and $\gamma$ then $\phi(w(g,\gamma))$ is the identity. By the assumption that $w(g,\gamma)$ is a relation in $g$ and $\gamma$ and the definition of $\phi$, we see that $\phi(w(g,\gamma))$ will be the identity if and only if $w$ has zero exponent sum in $\gamma$. Since we assumed that $\gamma^{n}\notin\langle\langle g\rangle\rangle_{K}$ for all nonzero $n$, it is immediate that any relation in $K$ will have zero exponent sum in $\gamma$.
It thus suffices to find the desired element $a\in C_{\Gamma}(g)$; we simply take a generator $a\in\mathbb{Z}$ of the infinite cyclic group in the free product. By construction, $(1,a)$ commutes with $(g_{1},1)\in K_{3}$, and the commutator $(1,a),(\gamma_{1},\gamma_{2})]$ is non-trivial and commutes with $(g_{1},1)$. ∎
@@
Let $G=\langle g,h\rangle$ be a non-cyclic and torsion-free group, and let $H$ be a free product with amalgamation of two copies $G_{1}=\langle g_{1},h_{1}\rangle$ and $G_{2}=\langle g_{2},h_{2}\rangle$ of $G$, wherein the subgroups $\langle g_{1}\rangle$ and $\langle g_{2}\rangle$ being amalgamated into a single cyclic group $\langle g\rangle$. Consider now the subgroup $K=\langle h_{2}^{-1}h_{1},g\rangle$. Then for all nonzero $n\in\mathbb{Z}$, we have
@@
We handle the case of $G$ cyclic, and then the case of $G$ non-cyclic. In the case where $G$ is cyclic, then it must be the case that for some $n\in\mathbb{Z}$ with $|n|>1$ we have that $h^{n}=g$. Indeed, if $g^{n}=h$ then $h$ must commute with $f$ which contradicts our assumption. We may produce the required overgroup by amalgamating a copy of $\mathbb{Z}^{3}$ generated by elements $\{a,b,c\}$ with $G$, by identifying the cyclic subgroups $\langle c\rangle$ and $\langle g\rangle$. Using the normal form for amalgamated free products, it is clear that $a,[b,h]]\neq 1$. Moreover,
@@
since $\alpha,g]=1$ for each $\alpha\in\{a,b,h\}$.
Assume now that $G$ is non-cyclic. Let $Q$ be the group $H$ obtained by amalgamating two copies of $G$ along their respective copies of $g$, and let $\beta$ be the automorphism of $Q$ which fixes $g$ and exchanges the two copies $h_{1}$ and $h_{2}$ of $h$ in $Q$, coming from the two amalgamated subgroups. We set $\Gamma_{0}$ to be the semidirect product of $Q$ with $\mathbb{Z}$, where $\mathbb{Z}$ acts on $Q$ by $\beta$. The group $Q$ is torsion-free by standard methods; see the discussion of free products with amalgamation in [14], for instance. Therefore $\Gamma_{0}$ is torsion-free as well.
We identify the generator of $\mathbb{Z}$ as an element $b$ of $\Gamma_{0}$, and we note that $b$ centralizes $g$. Observe that $b,h_{1}]=h_{2}^{-1}h_{1}$; we call this commutator $\gamma$, which clearly represents a non-trivial element of $\Gamma_{0}$.
Let $K=\langle g,\gamma\rangle$. By Lemma 3.3, we have that $\gamma^{n}\notin\langle\langle g\rangle\rangle_{K}$ for any $n\in\mathbb{Z}\setminus\{0\}$, and so by Lemma 3.2, the group $K$ embeds in a torsion-free overgroup $\tilde{K}$, wherein some non-trivial $a\in\tilde{K}$ centralizes $g$ and satisfies $a,\gamma]\neq 1$ and $a,\gamma]$ centralizes $g$. Now, amalgamate $\tilde{K}$ and $\Gamma_{0}$ over their respective copies of $K$, and call the resulting group $\Gamma$. In $\Gamma$, we have
@@
and all three of $a$ and $b$ and $a,[b,h_{1}]]$ centralize $g$. Since $\langle g,h_{1}\rangle$ is an isomorphic copy of $G$, we are done. ∎
We now need only prove Lemma 3.3. For this, we use Van Kampen diagrams over $H$; a reader unfamiliar with Van Kampen diagrams and disk diagrams may consult Chapter 5 of [9] or [11], for example. Of course, Van Kampen diagrams will generally depend on a particular choice of presentation. When we construct $H$, we have two copies of $G$, and as such there are two copies of the set of relations $\mathcal{R}_{0}$, which we call $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, respectively. Thus, relations in $\mathcal{R}_{i}$ involve $g=g_{1}=g_{2}$ and $h_{i}$ only. We fix the resulting presentation for $H$ once and for all.
@@
Suppose that in $G$ there are nonzero integers $n,m\in\mathbb{Z}$ such that $hg^{n}h^{-1}=g^{m}$. Then in the group $H$, we have $g^{n},h_{2}^{-1}h_{1}]=1$.